### v8

# Game-RL: Synthesizing Multimodal Verifiable Game Data to Boost VLMs’ General Reasoning

Abstract
Vision-language reinforcement learning (RL) has primarily focused on narrow domains (e.g. geometry or chart reasoning). This leaves broader training scenarios and resources underexplored, limiting the exploration and learning of Vision Language Models (VLMs) through RL. We find video games inherently provide rich visual elements and mechanics that are easy to verify. To fully use the multimodal and verifiable reward in video games, we propose Game-RL, constructing diverse game tasks for RL training to boost VLMs general reasoning ability. To obtain training data, we propose Code2Logic, a novel approach that adapts game code to synthesize game reasoning task data, thus obtaining the GameQA dataset of 30 games and 158 tasks with controllable difficulty gradation. Unexpectedly, RL training solely on GameQA enables multiple VLMs to achieve performance improvements across 7 diverse vision-language benchmarks, demonstrating the value of Game-RL for enhancing VLMs’ general reasoning. Furthermore, this suggests that video games may serve as valuable scenarios and resources to boost general reasoning abilities.
Abstract
[…]
1 Introduction
Vision language models have achieved impressive progress in basic tasks such as image description and vision question answering. However, they still struggle with diverse and complex tasks that require multi-step reasoning in real-world scenarios [18, 33]. One key reason is that vision-language RL primarily focused on narrow domains, mainly including geometry [22, 35] and chart reasoning [8], as shown in Table 1. This leaves broader training scenarios underexplored, limiting the exploration and learning of VLMs through RL [18, 14, 10, 36].
We recognize that video games inherently have three advantages: First, they can provide rich visual elements and scenes with texts. Second, their mechanics are easy to verify, so we can synthesize reliable tasks with verifiable results. Third, their environments are fully controllable and easy to modify, providing convenience for controlling the difficulty. Althrough existing works found game is a good arena to evaluate VLMs reasoning ability [33, 21, 32, 13], they did not apply training on it. One possible reason is they did not transform game data into Visual Question and Answer (VQA) format to adapt to training, as shown in Table 1.
To fully use the multimodal and verifiable rewards in game, we propose Game-RL, constructing game scenarios for RL training to boost VLMs general reasoning ability. Then we propose Code2Logic, a novel approach adapts game code to synthesize game reasoning task data, as shown in Figure 1. Code2Logic establishes mapping from game code to reasoning logic. Using Code2Logic, we have constructed the GameQA, which is a Visual question Answer (VQA) dataset to train and evaluate the reasoning capabilities of VLMs. GameQA includes 30 games, 158 tasks and 140K questions with controllable difficulty gradation, covering 4 cognitive categories, as shown in Figure 2 and Figure 3. Unexpectedly, despite training solely on game tasks, using Group Relative Policy Optimization (GRPO) [7], several SOTA open-source VLMs demonstrated out of domain generalization, specifically Qwen2.5-VL-7B achieving 2.33% improvement across 7 diverse vision-language benchmarks.
Table 1: Our GameQA dataset extends RL training scenarios for VLMs to the domain of video games, providing diverse verifiable game tasks along with controllable difficulty. Size means sum of the number of VQA pairs in train set and test set. Ration. Annot. means has annotated reasoning process. 3D Scene means some games of it are 3D scene. Detailed introduction can be found in Section 6.
Reasoning Related Work Train Game Ration. 3D Adjustable
Domain Set Count Annot. Scene Difficulty
MultiMath (Peng et al. [22]) N/A
Geo170k (Gao et al. [6]) N/A
MathV360K (Jiang et al. [11]) N/A
Game ING-VP (Zhang et al. [33]) 6
BALROG (Paglieri et al. [21]) 6
VideoGameBench 23
(Zhang et al. [32])
VCbench (Li et al. [13]) 10
GameQA (Ours) 30
Our main contributions are as follows:
  • We propose Game-RL, constructing game tasks for RL training to boost VLMs general reasoning ability. To obtain training data, We propose Code2Logic, a novel approach that adapts game code to synthesize diverse game reasoning task data.
  • Using Code2Logic, we developed the GameQA dataset, which includes 30 games, 158 tasks and 140K questions with controllable difficulty gradation.
  • Our experiments show that multiple VLMs achieve performance improvements across 7 diverse vision-language benchmarks through RL training solely on GameQA, demonstrating the value of Game-RL for enhancing VLMs’ general reasoning. This also suggests that video games may serve as valuable scenarios and resources to boost general reasoning abilities.
2 Code2Logic: Synthesizing Multimodal Verifiable Game Data
We propose Code2Logic to synthesize verifiable game reasoning tasks data via adapting game code. Code2Logic comprises three core steps: game code construction, task template design, and data engine construction for dataset generation, as shown in Figure 1. Code2Logic establishes a mapping from game code to reasoning logic.
Figure 1: Overview of Code2Logic approach. The process involves three main steps: (1) using LLMs to construct game code. (2) LLM-assisted design of the task templates including question and analysis templates based on the generated game code. Each task template condenses one type of reasoning pattern in the game. (3) Using LLMs to construct a data engine that directly reuses the core game code from the first step, including functions like move. (4) After these main steps, the data engine is executed to fill in the task templates developed in Step 2 and generate data samples, as illustrated in the Final Result Section.
2.1 Game code construction
The first step is to construct the code of the target video game. Game code defines the state space and contains some core functions encoding the transition rules of the game. These functions can be reused and adapted for data engine construction.
Taking Sokoban as an example, it is simple enough to be easily built with one-line prompt using LLMs such as Claude 3.5 and GPT-4o. Sokoban game state is only composed of wall, player, box and target, while action only includes moving in four directions. The prompt used for game code generation and pseudocode of the game is illustrated in Figure 1 (Step 1). Meanwhile, the Sokoban move function within Sokoban code can inspire the "State Prediction" questions, such as "Predict where will player end up after these steps", and this function can be reused for data engine construction.
2.2 Game task and QA Template design
The second step is to design game task and question-answer templates (QA Templates) according to each video game. Game task design base on the visual element and action space of video game generated in the first step. Taking Sokoban as an example, we can design a task about predicting the position of player after moves according to a game state screen, so we can propose a question and an answer. The question is "If the player moves left, up, down, down, where is the player?" and the answer is "The position of player is (2,2)." Then we can generalize these questions and answers into a QA Template, as shown in figure 1 (Step 2). Meanwhile, we can fill contents into a QA Template to get many instances. Detailed QA Templates of Sokoban is provided in Appendix 14.1.
In GameQA dataset, each task is defined as a group of samples from the same game that share the same QA template. We categorize the reasoning tasks into three types, including "Target Perception Task", "State Prediction Task" and "Strategy Optimization Task", with the detailed descriptions shown in Appendix 12.1.
For a task, we use LLMs to assist the design and refinement of the QA template, an example prompt for such refinement is illustrated in Figure 1 (Step 2). Additionally, LLMs can also design tasks, with example prompt shown in Appendix LABEL:app:supp_prompt. In summary, designing a QA template is equivalent to extracting one type of reasoning pattern from the video game code obtained in the first step. The third step then further instantiates these reasoning task templates into a dataset.
Figure 2: Four game examples from GameQA: 3D Reconstruction, Tangram, Sudoku, and Sokoban, each representing distinct cognitive categories. Each game displays two VQA examples consisting of: (a) current game state visualization, (b) a targeted question, and (c) step-by-step reasoning with the answer. GameQA transforms complex game-playing tasks into this structured VQA format. See Appendix LABEL:app:games_example_samples for more VQA examples of some representative games.
2.3 Data engine construction for data samples generation
The third step is data engine construction for samples generation. The data engine is a program that, when executed, automatically generates task instances in batch according to task templates. LLMs are used to assist in constructing the data engine’s code based on the game code obtained in the first step. The prompt for guiding LLMs to construct the data engine code is shown in Figure 1 (step 3). The pseudocode for the data engine is shown in Figure 1 (step 3). Taking Sokoban’s state prediction task as an example, the data engine program is constructed from four core modules:
  • Game environment initialization: This module randomly generates Sokoban environment by directly reusing the initialization functions from the Step 1’s game code. Sokoban environment includes the position of player, boxes, goals and walls.
  • Proposing task instance: This module proposes a state prediction task instance by generating a multi-step random movement sequence.
  • Solving task instance: This module generates a solution using a corresponding algorithm. For state prediction task, the algorithm simulates each action by reusing the movement logic from game code of Step 1, which handles all situations of movement including collisions and box pushing.
  • QA data construction: This module fills the task templates. The movement sequence will be filled into the question template to become a question instance. The state transition trajectory will be filled into the answer template to become an answer instance.
2.4 Data quality assessment and data augmentation
Table 2: GameQA train and test set statistics summary. The full table is in Appendix 13.1.
Statistic Category Train Set Test Set
Total Games 20 30
In-Domain 20 20
Out-of-Domain 0 10
Total Tasks 102 158
Total Questions 126,760 15,047
Multiple Choice 86,520 10,518
   Avg. Choices 7.10 7.05
Fill-in-the-Blank 40,240 4,529
Unique Images 74,620 8,620
We implemented manual quality verifications at each step of our method.
Game code verification: We verify the correctness of the game code generated by the models by manually running the game program. If errors are found, including vision issues in the game display or code logic errors, they are fed back to the LLM for regeneration. For complex game features that LLMs cannot generate correctly, we retrieved relevant open-source code and provided it to the LLM. Games created with external code are shown in Appendix 12.4.
Data engine validation: During the data engine development process, LLMs generate an initial version based on the prompt. This version is then manually tested. If the generated reasoning questions and answers do not conform to the templates from the second step or contain errors, the LLM will be instructed to regenerate the data engine.
Data augmentation and filtering: Once the data engine is constructed, data samples can be easily generated in batch, yet the answer processes often exhibit repetitive patterns. We therefore employed LLM paraphrasing to perform data augmentation, detailed in Appendix 13.3. Additionally, we applied data filtering to ensure the augmented samples were correct, of appropriate length, and without excessive textual repetition. Details of our data quality assurance process can be found in Appendix 13.4. Time spent on these main steps for each game in GameQA can be found in Appendix 12.3.
Figure 3: Overview of the GameQA dataset. The 30 games in GameQA can be classified into four categories based on the core abilities required to solve game tasks. Appendix 13.2 provides definitions of the four game categories. Games chose as Out-of-Domain are not used for training; instead, they are used to test the generalization performance after the model has been trained on In-Domain games.
3 The GameQA dataset
The GameQA dataset is a visual-language question answering dataset that transforms game-playing tasks into a Visual Question Answering (VQA) format, as shown in Figure 2. GameQA includes 30 games, 158 distinct tasks and approximately 140K questions in total.
Dataset composition: The dataset has 30 different games classified into four categories based on the core abilities required to solve game tasks: "3D Spatial Perception and Understanding", "Pattern Recognition and Matching", "Multi-step Reasoning", and "Strategic Planning", detailed descriptions shown in Appendix 13.2. The criteria on choosing these 30 games are listed in Appendix 12.2. As illustrated in Figure 3, the 30 games are divided into two sets: 20 In-Domain games for model training and 10 Out-of-Domain games for testing generalization performance. The Out-of-Domain set was held out during training.
Dataset question types: To ensure the final answers are verifiable, all questions in the dataset are either multiple-choice or fill-in-the-blank. The multiple-choice questions typically have 7-8 options, while the fill-in-the-blank questions require simple answers such as numbers or coordinates, with detailed statistics shown in Table 2.
Difficulty levels: Each task is assigned with one of three difficulty levels of questions ("QA Level"). And data samples are divided into 3 difficulty levels based on image complexity ("Plot Level"), namely game state or grid size, controlled by code parameters, providing three perception and reasoning difficulty levels for the tasks. Examples demonstrating these difficulty levels are in Appendix LABEL:app:games_example_samples. The training and evaluation result about difficulty are shown in Table 12 and Table 6.
4 Game-RL: RL on game tasks
4.1 RL algorithm
We use the Group Relative Policy Optimization (GRPO) [25] as our algorithm. We use the standard formulation from DeepSeek, with the loss function shown below and the hyperparameters shown in Appendix 10.3.2.
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}[q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)]\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}$
$\left\{\min\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]-\beta\mathbb{D}_{KL}[\pi_{\theta}||\pi_{\text{ref}}]\right\}$
4.2 Reward design
LLM as a judge to give reward. Given the high diversity of responses of our dataset, traditional rule-based judge methods (e.g. text matching) suffer from insufficient accuracy in answer matching. We chose the quantized version of Qwen2.5-32B-AWQ [23] as an evaluator model to assist in scoring the generated outputs.
Outcome reward design of RL The reward signal used for training was based solely on the correctness of the model’s final answer. Both the reference answer and the answer generated by the policy model are inputted into this evaluator model. The evaluator model then determines whether the generated answer matches the reference answer. The reward allocation adheres to a strict binary principle: a reward of 1 is assigned if they match, and a reward of 0 is assigned otherwise.
5 Experiment
Table 3: Evaluation results on general vision benchmarks. These results show that game-only training enhances general reasoning capability of VLMs. We fine-tuned four VLMs (InternVL2.5-8B [3], InternVL3-8B [4], Qwen2.5-VL-7B [2], and LLaVA-OneVision-7B [12]) on 5K GameQA samples using GRPO, resulting in the models Game-RL-InternVL2.5-8B, Game-RL-InternVL3-8B, Game-RL-Qwen2.5-VL-7B, and Game-RL-LLaVA-OV-7B. The percentage of performance improvements compared to the vanilla model is denoted by ($\uparrow$). Best performance per section is indicated in bold.
Models
Avg.
($\uparrow$)
MathVista MathVerse MMBench MMMU CharXiv MathVision MMMU-Pro
Baseline
Random 14.84 17.90 12.74 26.37 24.67 0.00 10.03 12.19
Proprietary Multimodal Large Language Models
GPT-4o 56.9 63.8 50.2 86.0 69.1 47.1 30.4 51.9
Claude-3.5-Sonnet 59.5 67.7 56.8 78.5 68.3 60.2 33.3 51.5
Gemini-2.5-Pro 73.1 77.7 65.9 88.3 79.7 62.9 66.0 71.2
Open-Source Multimodal Large Language Models
Qwen2.5-VL-32B 60.70 77.40 60.41 88.13 63.83 47.50 33.80 53.83
Ovis2-34B 57.91 71.50 53.71 88.73 60.91 49.10 35.93 45.48
InternVL2.5-38B 55.62 68.60 48.38 86.93 57.53 41.50 39.40 46.98
LLaVA-OV-72B 49.51 58.60 46.85 83.13 52.74 35.20 33.87 36.18
Qwen2.5-VL-72B 60.95 75.50 56.87 86.80 65.34 48.10 40.60 53.45
InternVL2.5-78B 57.96 70.20 52.34 88.33 61.84 42.70 39.93 50.38
InternVL2.5-8B 45.89 57.50 36.04 81.93 47.96 31.70 28.87 37.25
Game-RL-InternVL2.5-8B 47.91 (+2.02) 61.70 37.11 83.87 50.06 32.00 31.93 38.69
InternVL3-8B 54.48 69.10 50.10 86.00 57.88 39.10 35.33 43.84
Game-RL-InternVL3-8B 55.88 (+1.40) 73.00 50.71 86.20 58.34 39.90 37.93 45.10
Qwen2.5-VL-7B 49.94 66.80 45.08 83.67 49.01 37.70 30.80 36.49
Game-RL-Qwen2.5-VL-7B 52.27 (+2.33) 68.20 47.97 83.53 50.53 42.70 33.07 39.89
LLaVA-OV-7B 41.23 55.60 33.05 81.13 41.07 27.10 23.40 27.26
Game-RL-LLaVA-OV-7B 42.27 (+1.04) 58.20 34.92 82.53 41.31 27.30 23.07 28.58
Table 4: GameQA’s generalization is competitive compared to outstanding geometry visual reasoning datasets. Based on Qwen2.5-VL-7B, we applied the same GRPO method on 5k GameQA samples, 8k samples from MAVIS, 8k Multimodal-Open-R1 samples, 8k samples from MultiMath respectively, a mixture of 5k GameQA and 8k MultiMath samples, to conduct comparative training.
Models
Avg.
($\uparrow$)
Out of Domain Games
Avg.
($\uparrow$)
General Vision Benchmarks
3D Spatial
Perc. & Under.
Pattern Recog.
& Matching
Multi-step
Reasoning
Strategic
Planning
Math
Vista
Math
Verse
MMBench MMMU CharXiv
Math
Vision
MMMU-Pro
Qwen2.5-VL-7B 27.09 23.60 29.20 26.21 29.34 49.94 66.80 45.08 83.67 49.01 37.70 30.80 36.49
+ MAVIS (GRPO) 27.61 26.80 28.25 28.42 26.98 51.53 67.90 46.16 83.62 50.45 39.20 34.98 38.42
(+0.52) (+1.59)
+ Multimodal-Open-R1 28.33 24.87 27.86 29.93 30.64 51.86 67.63 48.09 83.78 49.78 40.20 34.89 38.65
(GRPO) (+1.24) (+1.92)
+ MultiMath (GRPO) 28.38 28.10 28.45 27.45 29.53 52.81 69.36 47.99 84.13 53.44 40.83 33.92 39.91
(+1.29) (+2.87)
+ GameQA (GRPO) 29.87 27.00 28.52 31.49 32.46 52.31 68.70 48.72 83.16 50.21 41.40 34.27 39.74
(+2.78) (+2.37)
GameQA,
MultiMath
30.93 28.20 28.28 34.26 32.99 53.23 69.20 48.02 84.73 53.21 42.10 34.47 40.89
(+3.84) (+3.29)
5.1 Experiment settings
The detailed description of data preparation, training processes, and evaluation method can be found in Appendix 10.
Hyperparamters of GRPO training We rollout 12 samples per questions. The model was trained for one epoch. The learning rate was 2e-7, with a 5% warm-up. The clipping value $\epsilon$ was set to 0.2 and the KL-divergence coefficient $\beta$ was set to 0.04.
Benchmarks We evaluated the models on a set of vison-language reasoning benchmarks consisting of our GameQA test set and public general benchmarks.
  • GameQA benchmark This test set includes around 500 question-answer pairs for each of 30 games, totaling 15,047 samples.
  • General benchmarks We used the MMMU validation set [30] for testing general multimodal understanding, and included MMMU-Pro [31], which features 10-option multiple-choice questions. To assess mathematical reasoning in visual contexts, we used MathVista [19] (testmini, 1,000 samples), MathVerse [34] (testmini, 3,940 samples), and MathVision [28] (open subset, 3,040 questions). For general visual understanding, we adopted MMBench [15] (validation set), and for chart-based reasoning, we used CharXiv [29].
5.2 Main results
By analyzing the models’ performance on in-domain tasks, related out-of-domain game tasks, and general vision-language benchmarks, we can find that:
Game-only training enhances general reasoning capability of VLMs Training on the GameQA dataset significantly improved performance on the GameQA test set. Moreover, models trained on GameQA exhibited strong generalization to unseen games, with accuracy improvements ranging from 1.16% to 3.82% (Table 5). On broader general vision benchmarks, these models showed robust generalization, achieving consistent performance gains across all general vision reasoning tasks. These results suggest that the models successfully learned transferable visual understanding and reasoning abilities from the GameQA dataset (see Table 3).
GameQA’s out-of-domain generalization is competitive compared to outstanding geometry visual reasoning datasets To better understand the GameQA dataset’s advantages, we compared it with the outstanding geometry visual reasoning datasets, include MAVIS [35], Multimodal-Open-R1 [16] and MultiMath [22]. MAVIS includes various geometry and function problems. Multimodal-Open-R1 is a geometry-centered dataset. MultiMath is a comprehensive and diverse multimodal mathematical dataset. The results (Table 4) show that despite having fewer training samples (5k vs. 8k) that are also out-of-domain for geometry and function tasks, the GameQA-trained model is competitive compared to its counterparts trained on geometry or function data, where general vision benchmarks would be considered in-domain. These results suggest that GameQA enables stronger out-of-domain generalization, even when using less data from a mismatched domain.
GameQA enhances model performance in mixed training. We trained Qwen2.5-VL-7B using GRPO on a mixture of 5k GameQA and 8k MultiMath samples, as shown in Table 4. This results suggest that GameQA can bring extra benefits when mix training with other dataset.
Figure 4: The scaling effect of training data quantity on general vision benchmarks for Qwen2.5-VL-7B-Instruct. The model was trained on a total of 20k samples (20 games) and evaluated every 1,000 samples. To clearly demonstrate the upward trend, the results are divided into three stages and presented using bin averaging (as described in Section 11.3).
Figure 5: The scaling effect of GameQA. As VLMs are trained on an increasing number of distinct games, their performance on general visual benchmarks improves. Game selection is shown in Table 7.
Figure 6: Impact of GRPO on visual perception performance on general visual benchmarks. Two pie charts and one example below illustrate how performance improves after GRPO.
5.3 Scaling effect of game samples on generalization
We trained the Qwen2.5-VL-7B model on a GameQA subset of 20,000 samples from 20 games using the GRPO method. As shown in Figure 4, the model’s performance score demonstrates a overall upward trend on 7 general vision benchmarks as the amount of training data increases. This indicates that scaling up training on GameQA data effectively enhances the VLM’s general reasoning abilities.
5.4 Scaling effect of game diversity on generalization
With around 5,000 total training samples, we trained the Qwen2.5-VL-3B on GameQA subsets with 4 and 20 distinct games (Table 7). The results (Figure 6) show a positive correlation between game diversity and generalization ability. This suggests that scaling game diversity makes better generalization, enabling the model to acquire more robust visual understanding and reasoning abilities.
5.5 Qualitative analysis
To confirm that GRPO substantially enhances visual perception and text reasoning abilities of models, we manually analyzed 790 cases randomly sampled from the results of InternVL2.5-8B, containing responses before and after GRPO. The results (Figure 6) confirm that after GRPO, the model demonstrates improved visual recognition of image elements and performs more precise reasoning. More statistics and cases are in Appendix 9.2.
In addition, our qualitative analysis of model performances across four game categories reveals common behaviors and challenges, detailed in Appendix LABEL:app_sec:gpt4o_case_study.
5.6 Advanced vision-language models perform notably worse than humans on the GameQA benchmark
Both leading open-source and proprietary models achieve average accuracy levels considerably lower than those of human (Table 5). This clear difference highlights the difficulty and high requirements of the GameQA benchmark, requiring not only accurate visual comprehension of game scenes but also the ability to carry out multi-step logical reasoning. The training and evaluation result about difficulty are shown in Table 12 and Table 6. Furthermore, our qualitative analysis and case study in Appendix LABEL:app_sec:gpt4o_case_study reveal that even the most advanced models currently struggle to match human-level understanding, particularly in tasks demanding deep reasoning. More experiments can be found in Appendix 11.
Table 5: Evaluation results on GameQA benchmark. The experimental setup is consistent with Table 3: four VLMs were fine-tuned on 5K GameQA samples using GRPO. We evaluate the models on both in-domain and out-of-domain game categories. The percentage improvement over the vanilla model is marked with a green upward arrow ($\uparrow$). The best performance for each category is presented in bold.
Models
Avg.
($\uparrow$)
In Domain Games
Avg.
($\uparrow$)
Out of Domain Games
3D Spatial
Perc. & Under.
Pattern Recog.
& Matching
Multi-step
Reasoning
Strategic
Planning
3D Spatial
Perc. & Under.
Pattern Recog.
& Matching
Multi-step
Reasoning
Strategic
Planning
Baselines
Human 84.75 85.18 80.74 84.46 88.62 81.61 79.17 73.81 81.27 92.19
Random 11.90 11.69 12.04 10.24 13.61 9.68 7.11 9.50 11.83 10.27
Proprietary Multimodal Large Language Models
GPT-4o 40.52 32.01 34.81 50.67 44.59 43.81 48.90 36.91 48.58 40.87
Claude-3.5-Sonnet 47.69 37.41 43.16 56.09 54.11 50.34 51.30 43.62 60.42 46.01
Claude-4-Sonnet 46.58 31.12 39.73 66.90 48.57 55.16 45.60 56.58 63.28 55.17
Gemini-2.5-Pro 58.95 46.93 52.79 74.62 61.46 67.60 57.60 77.37 77.62 57.80
Open-Source Multimodal Large Language Models
Ovis2-8B 24.98 19.92 24.43 27.21 28.37 26.99 29.70 20.70 34.14 23.41
InternVL3-9B 26.89 21.86 22.53 32.54 30.65 26.73 25.20 30.38 32.14 19.18
LLaMA3.2-11B-Vision 19.69 19.12 16.48 21.30 21.86 18.04 18.40 14.92 17.73 21.11
Qwen2.5-VL-32B 34.09 28.26 30.99 40.27 36.83 35.97 32.90 33.02 44.03 33.94
Ovis2-34B 34.53 27.92 32.72 39.46 38.03 35.29 35.50 34.20 38.71 32.75
InternVL2.5-38B 30.04 23.39 25.86 36.82 34.08 32.42 30.60 33.96 39.35 25.79
InternVL3-38B 35.23 28.33 31.76 41.89 38.96 38.69 33.30 43.62 50.09 27.75
LLaVA-OV-72B 24.87 19.92 24.88 27.72 26.95 28.32 26.80 23.52 32.87 30.11
Qwen2.5-VL-72B 37.63 29.47 32.85 45.76 42.42 39.22 35.90 37.38 46.86 36.75
InternVL2.5-78B 32.35 27.15 28.84 39.53 33.90 35.30 32.80 37.26 42.41 28.75
InternVL3-78B 38.00 33.15 33.03 46.60 39.20 39.74 34.90 40.70 50.95 32.43
InternVL2.5-8B 22.22 20.39 17.18 25.34 25.97 20.05 20.80 22.45 18.88 18.07
Game-RL-InternVL2.5-8B 29.44 (+7.22) 26.74 26.05 29.51 35.44 23.87 (+3.82) 25.00 25.12 24.91 20.45
InternVL3-8B 26.51 22.53 21.91 30.18 31.43 27.64 29.60 27.44 29.62 23.91
Game-RL-InternVL3-8B 33.09 (+6.58) 27.94 28.52 36.81 39.07 28.80 (+1.16) 29.20 25.31 34.59 26.09
Qwen2.5-VL-7B 25.78 22.58 21.92 25.21 33.40 27.09 23.60 29.20 26.21 29.34
Game-RL-Qwen2.5-VL-7B 32.12 (+6.34) 26.80 26.88 33.34 41.45 30.51 (+3.42) 27.10 31.56 31.24 32.13
LLaVA-OV-7B 21.79 18.84 19.69 21.24 27.38 20.39 21.00 20.19 20.13 20.27
Game-RL-LLaVA-OV-7B 33.49 (+11.70) 29.87 31.10 30.96 42.03 23.34 (+2.95) 27.20 20.05 23.55 22.57
6 Related work
Multimodal reasoning data construction Currently, the data construction methods are mainly divided into two categories: human expert supervision and automated synthesis. Peng et al. [22] and Lu et al. [17] collect visual reasoning problems from textbooks, Lu et al. [18] constructs datasets through labeling by STEM students, but they are limited by the scarcity of high-quality data sources and the high cost of manual annotation. Lu et al. [17], He et al. [8], Gao et al. [6], Shi et al. [26] uses expert models to generate reasoning processes, but the results are limited by the performance of the expert model. Trinh et al. [27] and Zhang et al. [35] synthesize geometric reasoning data through procedural methods, but these methods are often designed for specific domains and have high transfer costs. Table 1 provides a comparison of existing vision language reasoning datasets.
Using game data to enhance VLMs reasoning capabilities Game environments provide well-defined rules and mechanics that are easy to verify. However, existing work has not fully leveraged the potential of game environments in visual reasoning data construction. Reed et al. [24] trains a general agent by tokenizing game images and action sequences, but it is difficult to generalize on out-of-domain game tasks and this method relies on expensive expert trajectory data; Paglieri et al. [21], Zhang et al. [33, 32], Li et al. [13] all established gaming environments for Vision-Language Models, but these were used exclusively for evaluation purposes. These limitations indicate that how to effectively use game data to enhance the reasoning ability of visual language models remains a critical problem that needs to be addressed. Table 1 provides a comparison of existing game reasoning benchmarks.
7 Conclusion
To explore broader training scenarios and resources for vision-language RL, we propose Game-RL to construct game tasks for VLMs’ RL training. We also propose the novel Code2Logic approach adapting game code to synthesize diverse game reasoning task data, thus obtaining the GameQA dataset for Game-RL. Multiple VLMs trained through RL solely on GameQA achieved performance improvements across diverse general vision-language reasoning benchmarks. This not only demonstrates the value of Game-RL for enhancing VLMs’ general reasoning abilities, but also suggests that video games may serve as valuable scenarios and resources to boost VLMs’ general reasoning.
Appendix Contents
8 Limitations and future work
Using reasoning processes to conduct Supervised Fine-Tuning (SFT) has not achieved satisfactory out-of-domain generalization. Therefore, future work could explore methods to better leverage these reasoning processes to enhance the model’s general capabilities, such as employing reinforcement learning based on process supervision. In addition, GameQA currently involves single-turn game question answering. Future work could focus on developing training and evaluation methods for multi-turn interactions in gaming scenarios.
9 More analysis
9.1 GameQA evaluation results by difficulty
Table 6: Evaluation results on GameQA benchmark by difficulty. Scores are broken down by question difficulty (QA Level) and image complexity (Plot Level) within in-domain and out-of-domain game sets. The percentage of performance improvements compared to the vanilla model is denoted by ($\uparrow$). Best performance per section is indicated in bold.
Models
Avg.
($\uparrow$)
In Domain Games by Difficulty
Avg.
($\uparrow$)
Out of Domain Games by Difficulty
QA Level Plot Level QA Level Plot Level
Easy Medium Hard Easy Medium Hard Easy Medium Hard Easy Medium Hard
Proprietary Multimodal Large Language Models
GPT-4o 41.56 48.10 40.63 32.95 47.69 38.60 37.80 44.03 55.63 44.45 32.05 54.49 42.98 33.71
Claude-3.5-Sonnet 48.90 59.26 46.07 36.70 55.49 45.97 43.80 50.94 62.80 46.15 43.80 60.58 49.19 42.26
Claude-4-Sonnet 48.80 58.36 50.14 42.23 56.45 48.99 47.81 56.03 66.71 55.91 45.45 62.80 54.31 50.37
Gemini-2.5-Pro 60.62 67.43 62.36 56.21 67.48 62.42 57.72 67.68 74.47 67.56 61.10 74.23 66.43 61.90
Open-Source Multimodal Large Language Models
Ovis2-8B 25.56 33.67 22.80 20.21 28.30 26.12 23.69 27.37 32.41 28.87 20.90 33.96 25.56 22.06
InternVL3-9B 27.45 34.92 24.20 21.15 30.42 26.94 24.43 26.54 30.45 23.23 25.86 34.41 23.93 20.63
LLaMA3.2-11B-Vision 19.67 24.51 18.54 15.96 20.71 20.22 19.16 18.35 19.37 20.32 15.41 20.99 18.14 15.68
Qwen2.5-VL-32B 34.80 41.52 33.43 27.18 38.07 34.93 31.09 36.59 47.99 33.60 28.16 46.87 32.25 29.86
Ovis2-34B 35.32 46.36 31.81 25.11 39.57 34.70 31.70 35.40 46.86 30.02 29.22 42.95 33.76 28.87
InternVL2.5-38B 30.83 40.08 28.21 22.68 35.55 30.36 27.23 32.50 38.80 29.71 28.93 41.52 30.74 24.47
InternVL3-38B 36.07 44.05 31.73 30.60 40.70 36.47 30.20 38.78 50.00 36.14 30.17 49.03 35.62 30.86
LLaVA-OV-72B 25.49 35.50 21.21 17.57 27.81 24.66 23.87 28.98 38.80 23.83 24.20 34.07 27.00 25.46
Qwen2.5-VL-72B 38.58 47.61 36.70 27.38 43.50 37.67 33.43 39.77 52.90 37.90 28.51 49.15 37.13 32.28
InternVL2.5-78B 33.05 41.90 30.58 24.00 36.58 32.30 30.00 35.40 43.96 34.57 27.69 45.22 32.97 27.20
InternVL3-78B 38.64 47.64 35.60 30.47 43.91 37.34 34.57 40.17 52.37 35.84 32.23 50.34 39.36 29.93
InternVL2.5-8B 22.31 27.12 21.98 18.35 24.65 22.66 21.37 19.78 18.36 21.29 19.72 24.52 17.78 16.67
+ GameQA (SFT) 47.33 (+25.02) 56.59 45.36 39.79 53.73 45.50 44.76 26.10 (+6.32) 27.84 21.95 28.39 30.43 27.37 20.07
+ GameQA (GRPO) 29.52 (+7.21) 36.37 26.95 25.90 34.23 28.14 27.91 23.67 (+3.90) 24.41 22.92 23.67 28.44 22.85 19.33
InternVL3-8B 26.85 33.82 26.37 19.79 28.91 28.20 24.90 27.49 32.29 26.62 23.55 32.59 26.46 22.99
+ GameQA (SFT) 51.08 (+24.23) 63.08 47.66 43.22 58.93 48.69 48.74 27.35 (-0.14) 37.56 22.07 22.31 35.38 24.83 21.19
+ GameQA (GRPO) 33.53 (+6.68) 39.73 32.17 28.99 37.16 32.37 32.80 29.14 (+1.65) 34.36 29.41 23.67 34.81 27.31 24.85
Qwen2.5-VL-7B 26.02 36.25 23.11 17.75 28.29 25.69 25.67 27.25 31.87 25.83 24.03 33.73 25.38 22.12
+ GameQA (SFT) 49.23 (+23.21) 60.19 47.47 39.13 55.56 47.65 46.60 30.33 (+3.08) 42.00 25.35 23.55 36.86 29.29 24.29
+ GameQA (GRPO) 32.41 (+6.39) 42.51 30.25 23.96 35.05 32.40 31.79 30.77 (+3.52) 38.68 25.96 27.57 38.17 28.87 24.66
LLaVA-OV-7B 21.65 29.28 19.24 17.67 23.19 22.03 22.20 20.37 22.99 17.04 21.02 24.40 20.31 16.05
+ GameQA (SFT) 46.25 (+24.59) 54.62 46.81 36.70 51.65 44.55 45.03 23.39 (+3.02) 29.56 20.86 19.72 29.35 23.51 16.79
+ GameQA (GRPO) 33.61 (+11.95) 41.41 31.26 29.94 38.83 31.57 33.53 23.32 (+2.94) 30.15 18.25 21.43 28.90 23.51 17.04
The fine-grained difficulty gradation in the GameQA dataset enables a more systematic evaluation of the models. As shown in Table 6, when either QA Level or Plot Level increases, the models’ accuracy scores generally show a noticeable decrease.
Figure 7: Impact of GRPO on text reasoning performance, with two specific improvement examples.
Figure 8: Impact of GRPO on visual reasoning performance on GameQA.
9.2 Error types analysis
We manually checked the reasoning process of 650 GameQA cases randomly sampled from InternVL2.5-8B before and after GRPO to identify the error types. We categorize errors as 3 types: Visual Perception Error, Text Reasoning Error and both. The results are presented in Figure 9. Besides, in Figure 7 and Figure8 we supplement to Figure 6 with two examples of text reasoning improvement and another example of visual reasoning improvement after GRPO training.
Figure 9: InternVL2.5-8B error types before and after GRPO show that GRPO increases correct ratio and reduces visual and reasoning errors. Two cases show the two types of error in detail. Solid and dash lines connected to two cases means both are before GRPO.
10 Experiment details
10.1 Human and random baselines
We included two baselines for comparison: human and random.
  • Human Baseline: We selected approximately 20 questions from each of the 30 games, resulting in 623 questions. These were grouped into 30 sets and assigned to STEM undergraduates unfamiliar with the games. Each question was presented in a PowerPoint slide (Figure 10) using python-pptx [^1] https://python-pptx.readthedocs.io/en/latest/, and responses were collected via an online questionnaire. [^2] https://www.wjx.cn/
  • Random Baseline: This represents the lower performance bound, calculated as the expected score from random guessing on multiple-choice questions, with fill-in-the-blank tasks contributing zero.
Figure 10: Two example PowerPoint slides demonstrating how the student participants, during the human baseline evaluation, took the tests and might write or draw on the slides.
10.2 Data preparation
10.2.1 SFT data preparation
  • Game data preparation. First, we used our data engine to generate 5000 task instances from 20 in-domain games, obtaining 100000 samples in total. Then we performed data augmentation, as shown in Appendix 13.3. Finally, we filtered these samples, as shown in Appendix 13.4. 40000 samples were selected for SFT.
  • Other dataset preparation. We prepared the Geo-Multi dataset for comparing the data quality. Geo-Multi consists of 40,000 samples, with 20,000 randomly sampled from MultiMath300K [22] and another 20,000 from Geo170K [6].
10.2.2 Reinforcement Learning Data preparation
  • Game data preparation. We sampled 5000 samples from 20 in-domain games. This sample size is smaller than that used for SFT to balance performance and computational cost, as the GRPO training process is relatively resource-intensive.
  • Other dataset preparation. We sampled 4k examples from each of the geometry and function splits of MAVIS [35]. For multimodal-open-r1-8k-verified [16], we used all the samples.
Table 7: Selection of games for the game diversity scaling experiment across four cognitive categories.
Game Set 3D Spatial Perception and Understanding Pattern Recognition and Matching Multi-step Reasoning Strategic Planning
4 Games 3D Reconstruction Tangram Word-Search TicTacToe
10 Games 3D Maze Spider Solitaire Tents, 2D Turing Machine Sokoban, Space Invaders
20 Games Rubik’s Cube Freecell, Tetris, Zuma, Color Hue Langton’s Ant, Rhythm Game, Star Battle Maze, Ultra TicTacToe
To investigate the impact of game diversity on model generalization, we constructed subsets of 4, 10, and the full 20 in-domain training games. The selection process aimed to maintain representation across our four defined cognitive categories by randomly selecting games from each category to form the smaller subsets. Table 7 details the specific games included in each experimental set for this scaling analysis.
10.3 Training details
10.3.1 Models
We trained four VLMs, InternVL2.5-8B [3], InternVL3-8B [4], Qwen2.5-VL-7B [2], and LLaVA-OneVision-7B [12] on our data.
10.3.2 Training hyperparameters
LoRA-based supervised fine-tuning hyperparameters In this setup, Low-Rank Adaptation (LoRA) [9] was applied to all linear layers of the language model. We trained the model for one epoch. The rank was set to 16, with alpha set to 32. The learning rate was 5e-5, including a 3% warm-up period followed by a constant rate.
GRPO-based reinforcement learning In this setup, we conducted full-parameter fine-tuning of the language model while freezing the visual encoder and projection layers. We rollout 12 samples per questions. We trained the model for one epoch. The learning rate was 2e-7, with a 5% warm-up. The clipping value $\epsilon$ was set to 0.2 and the KL-divergence coefficient $\beta$ was set to 0.04. More hyperparameters are listed in Table 8.
Table 8: GRPO Hyperparameters.
Hyperparameter Value
Learning Rate 2e-7
Batch Size 3
KL-divergence Coefficient ($\beta$) 0.04
Number of Generations 12
Temperature 1.0
Top-p 0.85
Top-k 50
Optimizer AdamW
Warm-up Ratio 0.05
Weight Decay 0.1
10.3.3 Compute resources
LoRA-based supervised fine-tuning For the models in the 7-8 billion parameter range, this LoRA-based SFT training was conducted on a single A800 GPU and the training duration lasted around 15 hours or less for 40k samples (1 epoch).
GRPO-based reinforcement learning
  • 7-8 billion parameter model Training on 5k samples (1 epoch) required approximately 22 hours, utilizing five A800 GPUs, including resources for the deployment of the evaluator model.
  • 3 billion parameter model The training process lasted approximately 18 hours for 5k samples (1 epoch) on four A800 GPUs, also including the evaluator model’s deployment.
GPU usage Training a 7-8 billion parameter model with GRPO-based RL required approximately 22 hours, utilizing five A800 GPUs. This GPU allocation included resources for the deployment of the evaluator model. For a 3 billion parameter model, the training process lasted approximately 18 hours on four A800 GPUs, with this count also inclusive of the evaluator model’s deployment. For the 32 billion parameter model, the LoRA-based training took approximately 22 hours using eight H20 GPUs with 141GB memory each, while the reward model was deployed on four A100 GPUs.
10.4 Inference and evaluation
Besides trained models, we also evaluated proprietary large-scale models such as GPT-4o (20240806) [20] and Claude 3.5 Sonnet (20241022) [1], and open-source models that represent the current state of the art, including InternVL3-78B and Qwen2.5-VL-72B.
First, for evaluation on both GameQA and general vision benchmarks, the inference and evaluation configurations were unified across the original models and our trained models, detailed below.
For inference, the inference temperature parameter is set to 0.2, and the prompt is shown in Appendix LABEL:app_sec:prompt_for_inference. We evaluated the generated answers using the LLM-as-a-judge approach [37], with Qwen2.5-72B-AWQ acting as the evaluator and the prompt for evaluation shown in LABEL:app_sec:prompt_for_evaluation. To improve the reliability of the evaluation results, we introduced a series of engineering optimizations that ensured consistent and accurate assessments.
Second, We adopted different model evaluation strategies for different types of experiments to balance efficiency, accuracy, and the reliability of our conclusions, detailed below.
  • Selecting the model checkpoint with best performance on the validation set for evaluation This approach is employed for the GRPO-based training experiments shown in Table 3, 5, 6 and the SFT experiments detailed in Appendix 11.1. We saved 10 evenly spaced model checkpoints and evaluated the model that achieved the best performance on the validation set. The validation set was created by splitting 1% of the training data. It was found that the selected model was generally the final model in the training process. This approach efficiently and directly reflects the ultimate performance of our methods in practical scenarios.
  • Evaluating the top three model checkpoints based on the validation set performance and averaging the evaluation results This is employed for the comparative training shown in Table 4 and the analytic experiments of Section 5.4, Appendix LABEL:sec:more_results_on_diversity and Appendix 11.2. This approach was chosen because the performance differences are often subtle and susceptible to randomness in the training process. Averaging across multiple top checkpoints helps smooth out training noise, enhances the robustness and sensitivity of the evaluation, and more accurately reflects the differences under various experimental settings. This method also reduces the impact of chance results, making our experimental conclusions more reliable and representative.
11 More Experiments
11.1 SFT experiments
Table 9: Evaluation results on GameQA benchmark. In-domain and out-of-domain game category results are shown. The percentage of performance improvements compared to the vanilla model is denoted by ($\uparrow$). Best performance per section is indicated in bold.
Models
Avg.
($\uparrow$)
In Domain Games
Avg.
($\uparrow$)
Out of Domain Games
3D Spatial
Perc. & Under.
Pattern Recog.
& Matching
Multi-step
Reasoning
Strategic
Planning
3D Spatial
Perc. & Under.
Pattern Recog.
& Matching
Multi-step
Reasoning
Strategic
Planning
InternVL2.5-8B 22.22 20.39 17.18 25.34 25.97 20.05 20.80 22.45 18.88 18.07
+ GameQA (SFT) 46.73 (+24.51) 42.38 45.55 51.56 47.44 25.81 (+5.76) 24.40 24.69 25.90 28.25
+ GameQA (GRPO) 29.44 (+7.21) 26.74 26.05 29.51 35.44 23.87 (+3.82) 25.00 25.12 24.91 20.45
InternVL3-8B 26.51 22.53 21.91 30.18 31.43 27.64 29.60 27.44 29.62 23.91
+ GameQA (SFT) 50.72 (+24.21) 46.91 45.42 55.91 54.63 26.02 (-1.62) 24.20 14.80 32.03 33.05
+ GameQA (GRPO) 33.09 (+6.58) 27.94 28.52 36.81 39.07 28.80 (+1.15) 29.20 25.31 34.59 26.09
Qwen2.5-VL-7B 25.78 22.58 21.92 25.21 33.40 27.09 23.60 29.20 26.21 29.34
+ GameQA (SFT) 48.37 (+22.59) 42.05 42.82 58.66 49.96 29.27 (+2.18) 26.80 21.16 31.79 37.32
+ GameQA (GRPO) 32.12 (+6.34) 26.80 26.88 33.34 41.45 30.51 (+3.42) 27.10 31.56 31.24 32.13
LLaVA-OV-7B 21.79 18.84 19.69 21.24 27.38 20.39 21.00 20.19 20.13 20.27
+ GameQA (SFT) 45.40 (+23.61) 38.50 39.47 54.78 48.84 22.74 (+2.34) 22.00 17.17 25.16 26.63
+ GameQA (GRPO) 33.49 (+11.70) 29.87 31.10 30.96 42.03 23.34 (+2.95) 27.20 20.05 23.55 22.57
Table 10: Evaluation results on general vision benchmarks. The percentage of performance improvements compared to the vanilla model is denoted by ($\uparrow$). Best performance per section is indicated in bold.
Models
Avg.
($\uparrow$)
MathVista MathVerse MMBench MMMU CharXiv MathVision MMMU-Pro
InternVL2.5-8B 45.89 57.50 36.04 81.93 47.96 31.70 28.87 37.25
+ GameQA (SFT) 45.06 (-0.83) 58.10 35.79 82.87 48.07 31.20 22.80 36.56
+ Geo-Multi (SFT) 43.84(-2.05) 53.60 36.90 82.80 43.17 30.40 27.80 32.22
+ GameQA (GRPO) 47.91 (+2.02) 61.70 37.11 83.87 50.06 32.00 31.93 38.69
InternVL3-8B 54.48 69.10 50.10 86.00 57.88 39.10 35.33 43.84
+ GameQA (SFT) 49.66(-4.82) 63.20 43.30 84.53 53.44 32.90 29.60 40.64
+ Geo-Multi (SFT) 48.52 (-1.42) 62.60 40.96 82.87 48.54 37.80 29.80 37.06
+ GameQA (GRPO) 55.88 (+1.40) 73.00 50.71 86.20 58.34 39.90 37.93 45.10
Qwen2.5-VL-7B 49.94 66.80 45.08 83.67 49.01 37.70 30.80 36.49
+ GameQA (SFT) 47.26 (-2.72) 63.00 37.31 83.07 47.49 38.60 25.73 35.62
+ Geo-Multi (SFT) 48.52 (-1.42) 62.60 40.96 82.87 48.54 37.80 29.80 37.06
+ GameQA (GRPO) 52.27 (+2.33) 68.20 47.97 83.53 50.53 42.70 33.07 39.89
LLaVA-OV-7B 41.23 55.60 33.05 81.13 41.07 27.10 23.40 27.26
+ GameQA (SFT) 35.83 (-5.40) 45.90 25.99 80.27 32.21 20.50 20.47 25.44
+ Geo-Multi (SFT) 40.35 (-0.88) 52.50 31.57 82.40 41.31 25.90 22.60 26.19
+ GameQA (GRPO) 42.27 (+1.04) 58.20 34.92 82.53 41.31 27.30 23.07 28.58
As described in Appendix 10, we performed additional supervised fine-tuning (SFT) experiments on four models. The results of these experiments are presented in Tables 9 and 10, and we found that:
The GRPO demonstrates an advantage over SFT in out-of-domain generalization, though SFT yields strong in-domain gains. When evaluating model performance, the effects of SFT and GRPO training methods showed clear differences across domains. Specifically, SFT training with the GameQA dataset led to substantial in-domain improvements: the InternVL2.5-8B and Qwen2.5-VL-7B models improved their average accuracy by 24.51% and 22.59% respectively on a test set covering 20 games, reaching final scores of 46.73% and 48.37% (Table 9). However, while SFT excels at improving in-domain performance, it can sometimes lead to performance degradation on general tasks as seen in 9, a phenomenon known as "catastrophic forgetting". Crucially, for training sets that belong to the mathematics domain, such as Geo-Multi (see Appendix 10.2), models trained on them can still exhibit a decline in general capabilities. This may reflect the tendency of the SFT method to overfit the model to specific domain data [5].
In contrast, when training on GameQA using the GRPO, models not only successfully avoided performance degradation on general tasks but also generally achieved performance improvements across multiple general visual benchmarks (Table 3). A typical example is the Qwen2.5-VL-7B model, which, after GRPO training, showed enhanced performance on challenging benchmarks such as MathVista, MathVerse, and CharXiv, reaching a level comparable to larger-scale models like InternVL2.5-38B.
Pure RL outperforms SFT-then-RL in out-of-domain generalization. To validate our choice of directly applying RL, we compare our approach to a two-stage SFT-then-RL pipeline on Qwen2.5-VL-7B. As shown in Table 11, while the two-stage pipeline yields strong performance on in-domain tasks, it leads to a significant performance drop of -2.5% on general vision benchmarks. This suggests that full-parameter SFT on a narrow, specialized domain like GameQA can cause catastrophic forgetting. In contrast, pure RL better preserves and enhances out-of-domain generalization capabilities, achieving a +2.33% improvement.
Table 11: Comparison of training pipelines on Qwen2.5-VL-7B. SFT is performed on 20k samples, followed by GRPO on 5k. Performance on general benchmarks is shown as a relative change from the baseline.
Training Stage In-domain Out-of-domain General Benchmarks
Games Games (Avg. Change)
Baseline (Before SFT) 25.89 26.92 0.0%
SFT 56.21 30.74 -3.9%
SFT-then-RL 58.08 31.96 -2.5%
Pure RL (Our Method) 32.12 30.51 +2.33%
11.2 Training data difficulty experiment
To analyze how reasoning difficulty impacts generalization, we categorized tasks into Easy (55-85% baseline accuracy), Medium (30-55%), and Hard (5-30%). We trained Qwen2.5-VL-7B on 5k samples from different difficulty combinations. Table 12 shows that training on a diverse mix of difficulties yields the best generalization. The model trained on "Easy+Medium+Hard" samples achieves the highest average score (52.74), outperforming models trained on simpler subsets. This confirms that the complex reasoning patterns in GameQA are crucial for improving general reasoning.
Table 12: Impact of training data difficulty on generalization. Results are averaged over the top three checkpoints. The baseline score is 49.94.
Training Data (5k) Average MathVista MathVerse MMBench MMMU CharXiv MathVision MMMU-Pro
Baseline 49.94 66.80 45.08 83.67 49.01 37.70 30.80 36.49
Easy 52.40 67.87 47.93 83.91 51.58 41.77 33.73 40.03
Medium 51.95 67.30 48.10 83.64 50.53 40.83 33.96 39.30
Hard 52.07 67.57 47.53 83.58 51.03 40.97 34.58 39.26
Easy+Medium 52.45 68.60 48.07 83.42 51.30 41.40 34.86 39.49
Easy+Medium+Hard 52.74 68.33 48.68 83.93 52.36 40.73 34.31 40.81
11.3 Data scale experiment
We extended our data scaling experiment up to 20k samples. While individual checkpoints exhibit fluctuations common in RL, a "binned averaging" analysis (Table 13) reveals a clear, monotonic positive trend. By averaging performance across training stages, we smooth out short-term noise and observe a stable improvement trajectory. The average performance on general benchmarks steadily increases from +2.31 in the early stage to +3.02 in the late stage, confirming that model performance continues to improve with more data from GameQA.
Table 13: Binned averaging analysis of data scaling effect on Qwen2.5-VL-7B. Performance gains over baseline are shown in parentheses.
Training Stage Sample Range General Vision Benchmarks In-domain Games
Baseline Model 0k 49.94 25.78
Early Stage Training 1k–5k 52.25 (+2.31) 27.68 (+1.90)
Mid Stage Training 6k–15k 52.57 (+2.63) 34.97 (+9.19)
Late Stage Training 16k–20k 52.96 (+3.02) 38.69 (+12.91)
12 Approach details
12.1 Task category
Target Perception Task focuses on visual perception and basic state awareness. State Prediction Task, building directly on the perceptual capacities, needs predictions about state transitions. Strategy Optimization Task then needs both perceptual and predictive capacities to find optimal solutions. This progressive structure helps organize reasoning skills from simple to complex.
The conceptual outline for each category, using Sokoban as an example, is as follows:
  • Target Perception Task: Queries static information within the game state. For instance, questions ask about the position or number of boxes, and the answers list the specific positions of each box by directly inspecting the current state.
  • State Prediction Task: Infers state changes following actions. For instance, questions predict the player’s position after a sequence of moves. Answers are derived by analyzing the initial state, simulating the execution of each step, recording the resulting state changes, and thus determining the final player position.
  • Strategy Optimization Task: Aims to find optimal solutions. For instance, questions ask for the shortest path to push a specific box to its designated target. The answer is derived by first analyzing the initial state to determine the optimal route for moving the box to its target, and then simulating the execution of this optimal action sequence.
12.2 Selection criteria of 30 games in GameQA
We selected these 30 games based on the following criteria.
  • Ability coverage The games need to cover a diverse range of reasoning abilities, including 3D Spatial Perception and Understanding, Pattern Recognition and Matching, Multi-step Reasoning, and Strategic Planning.
  • Code simplicity The code should be easy to construct, meaning they are simple enough to be programmed by an LLM, or are open-sourced.
  • Static game They should be static or can be transformed into a static state, so that problems can be solved from a static image.
12.3 Time spent on the main steps of Code2Logic for each game
Figure 11 illustrates the estimated time spent on implementing the main steps of the Code2Logic approach across all the 30 games in the GameQA dataset. The time ranges from a minimum of 4 hours to a maximum of 12 hours, with an average of 7.5 hours per game. Given that all annotators were undergraduates and that ten of them completed only a single game, the time required for a more experienced annotator could be significantly less than this average.
This average time investment is relatively cost-effective and appears highly acceptable, especially considering that once the code data engine is built, it can generate an unlimited number of data samples for training and evaluation purposes.
Figure 11: Estimated time (in hours) spent on implementing the main steps of Code2Logic across different games in the GameQA dataset, with an overall average of 7.5 hours per game.
12.4 Games generated using external open-source code
This section lists the games whose code is generated based on open-source code, as referenced in Section 2.1.
Spider Solitaire (Open-source code URL: https://github.com/rdasxy/spider_solitaire)
Based on an original Python implementation of Spider Solitaire, our code reused its core rules and GUI. And we simplified the game to a single suit (Spades) to reduce complexity, enriched initial setups through LLM-implemented random deals, and adapted the original game rules into detailed instructions.
Klondike (Open-source code URL: https://github.com/milorb/klondike)
Based on the original open-source Klondike Solitaire project built on Pygame [^3] https://www.pygame.org/news, we adapted the code using an LLM into an automated dataset generation tool. It reused the core game engine and introduced a method for generating diverse random initializations.
Space Invaders (Open-source code URL: https://github.com/leerob/space-invaders)
Based on the original Space Invaders game built with Pygame, we utilized its core elements and visual assets to generate static game scenes. Using an LLM, we converted the dynamic game into static scene snapshots for dataset generation.
13 More information about the GameQA Dataset
13.1 The GameQA dataset statistics
The full GameQA train and test set statistics table is shown in Table 14.
13.2 Definition of each category
3D Spatial Perception and Reasoning Game of this type involves the ability to perceive, plan, and reason in 3D space to complete tasks such as navigation and spatial transformation. For example, in 3D Reconstruction, the task is to reconstruct the stacking arrangement of 3D cubes from a side view. Solving this task requires 3D spatial reasoning to establish the relationship between the 2D view and the 3D view.
Pattern Recognition and Matching Game of this type requires capability on discerning and matching visual patterns related to object shapes, colors, combinations, and other regularities. For example, in Tangram, the task is to identify which piece can fill the empty space. Solving this task requires recognizing the shape of void and matching with given pieces.
Table 14: GameQA train and test set statistics summary. All lengths are calculated by words.
Statistic Category Train Set Test Set
Overall Counts
Total Games 20 30
Total Tasks 102 158
Total Questions 126,760 15,047
Image Statistics
Unique Images 74,620 8,620
Avg. Image Width (px) 511.00 504.10
Avg. Image Height (px) 475.73 468.98
Question Characteristics
Avg. Question Length 275.27 272.43
Avg. Analysis Length 106.85 144.89
- After Augmentation 300.79 -
Multiple Choice Questions 86,520 10,518
Avg. Choices for MCQs 7.10 7.05
Fill-in-the-Blank 40,240 4,529
Multi-step Reasoning Game of this type features multi-step reasoning and iteratively applying rules to reach the solution. For example, in Sudoku, the task is to infer which color should fill the empty space to ensure that no colors are repeated in the same row, column, or 3x3 grid. The "no repetition" rule needs to be applied repeatedly to deduct the correct color of a cell.
Strategic Planning Game of this type requires planning the optimal solution in optimization problems. For example, in Sokoban, the task is to plan the shortest path for pushing a box from the starting point to the target location.
13.3 Data augmentation
To prevent model overfitting to specific reasoning patterns, we employed an LLM-based reasoning paraphrase strategy using InternVL2.5-78B. Based on initial experiments showing that visual input could lead to errors due to the model’s insufficient visual capabilities, we provided it only with textual information, namely the question, original answer process and final answer, to rewrite the answer process. This approach enriches linguistic style and logical expression diversity while maintaining semantic consistency.
13.4 Data quality assurance
To ensure the high quality and reliability of our synthetic dataset, we implemented a quality assurance process, consisting of four stages:
  1. 1.
    Human Inspection: STEM students inspected initial samples to ensure logical correctness, clarity and completeness of questions, images and reasoning steps.
  2. 2.
    LLMs Check: Fed samples to GPT-4o and Claude-3.5 Sonnet to ensure model comprehensibility, identifying necessary refinements in the samples.
  3. 3.
    Post-Augmentation Verification: Manually verified reasoning accuracy in a random subset of augmented data.
  4. 4.
    Automated Data Filtering: Removed samples based on length, high repetition (>70% 4-gram overlap), or wrong answers, reducing the set from 1̃50k to 126,760.
14 Details on Sokoban task systhesis
14.1 Sokoban QA templates
Here we provide templates used for generating Sokoban puzzle samples, including the three problem types: Target Perception, State Prediction, and Strategic Optimization.
Target perception QA template The Target Perception template (Table 14.1) is used to generate questions about identifying the current position of game elements.
Table 15: Target perception QA template